A partnership has the following capital balances:

Henry (60% of gains and losses) $ 76,000
Thomas (25%) 90,000
Catherine (15%) 130,000

Anne is going to invest $75,000 into the business to acquire a 30 percent ownership interest. Goodwill is to be recorded. What will be Anne’s beginning capital balance?

Multiple Choice

  • $126,857

  • $88,800

  • $111,300

  • $75,000
Answer #1

Anne's Capital : 30% of ( Existing capital + Annes Capital)

$75000+ Goodwill = 0.3 ($296000+$75000) = $75000+ Goodwill= $111300

Goodwill ; 111300-75000 =36300

0.7 goodwill ;36300/.7 =51857 (Good will)

Annes capital balance = 75000+51857=$126857 (First option correct)